OpenStudy (anonymous):

*Medal for best answer* What are the x-intercept(s) of the graph of y + 20 = x2 − x? (−4, 0) and (−5, 0) (−4, 0) and (5, 0) (4, 0) and (−5, 0) (4, 0) and (5, 0)

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (ddcamp):

x intercepts are at y=0: 0 + 20 = x² - x x² - x = 20 x² - x - 20 = 0 (x+a)(x+b) = 0 Solve the system: a + b = -1 a * b = -20